### v1

# CoMPaSS: Enhancing Spatial Understanding in Text-to-Image Diffusion Models

Abstract
Text-to-image (T2I) diffusion models excel at generating photorealistic images, but commonly struggle to render accurate spatial relationships described in text prompts. We identify two core issues underlying this common failure: 1) the ambiguous nature of spatial-related data in existing datasets, and 2) the inability of current text encoders to accurately interpret the spatial semantics of input descriptions. We address these issues with CoMPaSS, a versatile training framework that enhances spatial understanding of any T2I diffusion model. CoMPaSS solves the ambiguity of spatial-related data with the Spatial Constraints-Oriented Pairing (SCOP) data engine, which curates spatially-accurate training data through a set of principled spatial constraints. To better exploit the curated high-quality spatial priors, CoMPaSS further introduces a Token ENcoding ORdering (TENOR) module to allow better exploitation of high-quality spatial priors, effectively compensating for the shortcoming of text encoders. Extensive experiments on four popular open-weight T2I diffusion models covering both UNet- and MMDiT-based architectures demonstrate the effectiveness of CoMPaSS by setting new state-of-the-arts with substantial relative gains across well-known benchmarks on spatial relationships generation, including VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%). Code will be available at https://github.com/blurgyy/CoMPaSS.
1 Introduction
Recent advancements in text-to-image (T2I) diffusion models [saharia_photorealistic_2022, ramesh_hierarchical_2022, rombach_high-resolution_2022] have transformed visual content creation and significantly shaped the modern digital life [betker_improving_nodate, MidJourney, RunwayML]. These models are capable of synthesizing photorealistic images with remarkable details [dhariwal_diffusion_2021], and are now widely applied across a range of creative and practical tasks [hertz_prompt--prompt_2022, lugmayr_repaint_2022, ruiz_dreambooth_2023, poole_dreamfusion_2022]. Despite these advancements, however, a key limitation persists: today’s diffusion models frequently fail to correctly render spatial relationships described in text. For example, when given seemingly simple spatial configurations like “a motorcycle to the right of a bear”, or “a bird below a skateboard”, models that excel in realism often struggle to accurately capture these spatial relationships (Fig. 1 left).
Figure 1: CoMPaSS enables existing T2I diffusion models to generate spatial configurations that are unlikely to occur naturally and unseen during training by improving their spatial understanding.
Given that T2I diffusion models have shown significant capability in encoding abstract attributes such as aesthetic style and quality [dhariwal_diffusion_2021, rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024], their consistent failure in handling spatial relationships warrants closer examinations. We consider that the limitation stems from how spatial priors are organized within the training data, and how these priors are represented to the diffusion model.
To better understand this limitation, we conduct a thorough investigation covering the data characteristics during training, and the language representation during generation. Our analysis identified two core issues underlying this spatial understanding deficiency. First, the organization of spatial related data in widely-used image-text datasets including LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022], CC-12m [changpinyo_conceptual_2021], and COCO [lin_microsoft_2014] is severely flawed. As shown in Fig. 2, spatial descriptions appear frequently ambiguous due to various reasons, such as inconsistent frame of reference, or non-spatial usage of spatial terms. This inevitably leads to the spatial terms being interpreted by the diffusion model as ambiguous. Second, commonly used text encoders including CLIP [radford_learning_2021, cherti_reproducible_2023] and T5 [raffel_exploring_2020] consistently fail to preserve spatial relationships in their embeddings. Through a proxy task where the text encoders are extensively tested to retrieve the spatial equivalent from three prompt variations (Tab. 1), we show that current text encoders struggle to interpret the accurate meanings of spatial descriptions. Furthermore, the two issues compound each other. Even if a text encoder could reliably capture spatial relationships, the lack of consistent and accurate spatial data would still limit model performance. Conversely, even high-quality spatial descriptions would be undermined by the encoder’s inability to interpret them accurately.
To this end, we seek to enhance the spatial understanding of text-to-image generative diffusion models with a comprehensive solution, which we call CoMPaSS (Comprehensive Method for Positional and Spatial Synthesis). Central to our approach is the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. By enforcing strict criteria for visual prominence, positional clarity, and size balance, SCOP identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. When applied to the COCO [lin_microsoft_2014] training split, SCOP yields a curated set of over 28,000 object pairs, each coupled with spatially-accurate text descriptions of the image content. To better exploit the accurate spatial priors, we additionally introduce Token ENcoding ORdering (TENOR), a plug-and-play module that enhances the preservation of spatial semantics from text input with explicit token ordering information, effectively compensating for the spatial understanding limitations of current text encoders. TENOR is compatible with any diffusion architecture, adds no extra parameters, and inflicts practically zero inference-time overhead.
Applying CoMPaSS to popular open-weight T2I diffusion models like SD1.4, SD1.5, SD2.1 and FLUX.1 sets new state-of-the-arts with substantial relative gains on well-known benchmarks, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%). In summary, our contributions are three-fold:
  • A comprehensive training framework, CoMPaSS, that significantly enhances spatial understanding in T2I diffusion models.
  • A systematic data engine, SCOP, that enforces principled constraints to identify and validate unambiguous spatial relationships between object pairs in images, enabling the curation of high-quality spatial training data.
  • A parameter-free module, TENOR, that further improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds negligible computational overhead.
2 Related Works
Text-to-image Diffusion Models.
Diffusion models have achieved very high-quality and diverse image synthesis in recent years [ho_denoising_2020, rombach_high-resolution_2022, dhariwal_diffusion_2021, peebles_scalable_2023]. To allow natural user input through human language, text-conditional diffusion models adopt datasets with large-scale image-text pairs [changpinyo_conceptual_2021, schuhmann_laion-400m_2021, schuhmann_laion-5b_2022, liang_rich_2024], with pre-trained text encoders to bridge the gap between the two modalities. GLIDE [nichol_glide_2022], DALLE-2 [ramesh_hierarchical_2022], LDM [rombach_high-resolution_2022], and the Stable-Diffusion family of models [rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024] employ CLIP [radford_learning_2021, cherti_reproducible_2023] for language conditioning. PixArt-$\alpha$ [chen_pixart-alpha_2023], SD3/3.5 [esser_scaling_2024], and FLUX.1 adopt the pure-language T5 [raffel_exploring_2020] model as their text encoders. Benchmarking the alignment of image content w.r.t. the input text description has been well explored [hu_tifa_2023, ghosh_geneval_2023, gokhale_benchmarking_2023, huang_t2i-compbench_2023-1, jayasumana_rethinking_2024, hu_ella_2024]. More recently, CoMat [jiang_comat_2024] leverages image-to-text (I2T) models [li_blip_2022, wang_git_2022, liu_visual_2023] to identify the overlooked condition information in generated images during training. ParaDiffusion [wu_paragraph--image_2023] and ELLA [hu_ella_2024] explore the use of large language models (LLMs) in T2I diffusion models.
Instead of trying to enhance the language understanding ability by using more advanced text encoders, we seek an orthogonal direction of facilitating text-image alignment with a data engine that ensures accurate spatial alignment between image and text in training data.
Spatial Control in Diffusion Models.
Controllable image generation has become a popular research topic recently. Several lines of works, including both training-based and test-time optimization methods have been explored to enhance the spatial controllability of generation. Previous works add image as conditioning input [zhang_adding_2023, mou_t2i-adapter_2023, ye_ip-adapter_2023] to enable layout control, but add computational overhead to both training and inference processes. Attention-based methods build upon the relation between the text-image attention map and the layout of generated image [hertz_prompt--prompt_2022], and improve compositional generation by explicitly supervising the attention maps either during training [wang_tokencompose_2024] or at inference-time [chefer_attend-and-excite_2023, chen_training-free_2023, phung_grounded_2024]. However, supervising attention requires explicit target signal for optimization. Although works like [phung_grounded_2024, feng_layoutgpt_2024] automate the generation of target signals with LLMs, optimization on attention maps still inflicts substantial memory and speed penalty. SPRIGHT [chatterjee_getting_2024] avoids test-time optimizations by fine-tuning a diffusion model on datasets recaptioned with a vision-language model [liu_visual_2023], but requires joint training of the text encoder in addition to the diffusion network, which still adds considerable training overhead.
In contrast to existing works, we propose a network module that improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds no extra parameters and inflicts practically zero computational overheads to either training or inference processes.
3 Approach
We present CoMPaSS, a comprehensive solution for improving spatial understanding in text-to-image diffusion models. In Sec. 3.1, we develop the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. In Sec. 3.2, we introduce the Token ENcoding ORdering (TENOR) module, a plug-and-play module that enhances the preservation of spatial semantics from text input by incorporating explicit token ordering information.
These components together form a comprehensive approach to improving spatial understanding in text-to-image diffusion models: SCOP provides high-quality spatial priors through carefully curated spatial relationships, while TENOR enables the model to effectively interpret and exploit these spatial priors during generation.
3.1 The SCOP Data Engine
High-quality spatial understanding in generative models requires training data with unambiguous spatial relationships. However, existing text-image datasets often contain problematic spatial descriptions. As shown in Fig. 2, 1) terms like “left” and “right” suffer from perspective ambiguity, referring inconsistently to viewer perspective or object-intrinsic orientation; 2) directional terms are frequently used in non-spatial contexts (e.g.the right choice”); 3) spatial relationships are often described with missing or incorrect reference objects, making the intended spatial configuration impossible to determine. These issues create significant challenges for models attempting to learn reliable spatial relationships from such data.
We address this limitation through the Spatial Constraints-Oriented Pairing (SCOP) data engine, which identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. As illustrated in Fig. 3, SCOP processes images through three key stages:
((a)) COCO [lin_microsoft_2014]
((b)) LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022]
((c)) CC-12M [changpinyo_conceptual_2021]
Figure 2: Examples highlighting common ambiguities and errors in spatial language annotations from COCO, LAION, and CC-12M datasets. (a1, a2) inconsistent frame of reference; (b1) non-spatial usage of spatial terms; (b2, c1, c2) missing or incorrect reference objects.
Figure 3: Overview of the Spatial Constraints-Oriented Pairing (SCOP) data engine. SCOP first (1) reasons through all possible pairs of objects in image, then (2) identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. Finally, it (3) decodes the extracted unambiguous object pair descriptors into images with accurate spatial descriptions
1. Relationship Reasoning.
The engine begins by identifying all possible pairwise relationships within each image. For every image $\mathbf{I}$, we first identify all valid object instances $\{O_{1},\cdots,O_{n}\}$, where the $i$-th object $O_{i}$ is associated with a bounding box $\mathbf{B}_{i}$ and its corresponding category label $\text{Category}(\mathbf{B}_{i})$. We then enumerate all possible pairs between the identified objects, creating a total of $(^{n}_{2})$ candidate pairs with their respective spatial relationships to evaluate.
2. Spatial Constraints Enforcement.
To ensure unambiguous spatial relationships between object pairs in images, we formalize a set of necessary constraints that collectively guarantee visual clarity and semantic validity. For any candidate object pair $(O_{i},O_{j})$, we evaluate five constraints:
  • Visual Significance: Objects must occupy a sufficient portion of the image to ensure their spatial relationship is meaningful:
    $\frac{\text{Area}(\mathbf{B}_{i}\cup\mathbf{B}_{j})}{\text{Area}(\mathbf{I})}>\tau_{\text{v}}.$ (1)
    A relatively small value of $\tau_{\text{v}}$ ensures the reasoned spatial relationship from last stage is a dominant feature of the image, rather than an incidental arrangement of background elements.
  • Semantic Distinction: Objects must belong to different categories, i.e.:
    $\text{Category}(\textbf{B}_{i})\neq\text{Category}(\textbf{B}_{j}).$ (2)
    This constraint eliminates ambiguous cases where multiple instances of the same object category could create reference confusion in spatial descriptions.
  • Spatial Clarity: Objects must be within a reasonable proximity for clear spatial reference. Formally, let $d(i,j)=\left\lVert\text{centroid}(\mathbf{B}_{i})-\text{centroid}(\mathbf{B}_{j})\right\rVert_{2}$, and $l(i,j)=\min\left(\text{diagonal}(\mathbf{B}_{i}),\text{diagonal}(\mathbf{B}_{j})\right)$, an object pair $(O_{i},O_{j})$ is said to satisfy spatial clarity when:
    $\frac{d(i,j)}{l(i,j)}<\tau_{\text{u}},$ (3)
    where $\tau_{\text{u}}$ upper-bounds the relative distance between $O_{i}$ and $O_{j}$, avoiding ambiguous spatial relationships between far-apart objects.
  • Minimal Overlap: Objects must maintain sufficient visual separation:
    $\frac{\text{Area}(\mathbf{B}_{i}\cap\mathbf{B}_{j})}{\min\left(\text{Area}(\mathbf{B}_{i}),\text{Area}(\mathbf{B}_{j})\right)}<\tau_{\text{o}}.$ (4)
    The overlap threshold $\tau_{\text{o}}$ preserves individual object visibility, while allowing natural spatial configurations like “cup on table” where partial overlap is expected.
  • Size Balance: The objects in a pair should have comparable visual prominence:
    $\frac{\min\left(\text{Area}(\mathbf{B}_{i}),\text{Area}(\mathbf{B}_{j})\right)}{\max\left(\text{Area}(\mathbf{B}_{i}),\text{Area}(\mathbf{B}_{j})\right)}>\tau_{\text{s}}.$ (5)
    With this constraint, both objects should contribute similarly to the spatial relationship, preventing cases where one object is too small to serve as a reliable spatial reference.
These constraints effectively filter out ambiguous spatial relationships while preserving natural object interactions, creating a strong foundation for clear and consistent spatial descriptions.
3. Relationship Decoding.
For object pairs satisfying all constraints, we adopt a representation system that encodes the spatial relationships with structured descriptors, rather than fixed text. Such descriptors can be decoded into image-text pairs at training time. For example, a descriptor $D=$(cup, $\mathbf{B}_{\mathtt{cup}}$) <above> (couch, $\mathbf{B}_{\mathtt{couch}}$)” can be decoded into $(\mathbf{I}_{\text{roi}},\mathbf{T})$, where $\mathbf{I}_{\text{roi}}$ is a crop of the original image that contains both $\mathbf{B}_{\mathtt{cup}}$ and $\mathbf{B}_{\mathtt{couch}}$, and $\mathbf{T}$ is a text prompt formatted from a predefined pool of templates, e.g.a cup on top of a couch”, or “an image of a couch below a cup”, etc. Such a decoding process captures the exact positions of the referred objects as well as their reasoned spatial relation, allowing for both accurate relation descriptions and flexible training signals.
The SCOP Dataset.
We automate SCOP on the COCO [lin_microsoft_2014] training split to curate a dataset specifically targeted at training T2I models with improved spatial understanding capabilities, covering over 28,000 object pairs across 15,000 images, with well-defined spatial relationships and accurate spatial descriptors. This dataset significantly improves the spatial understanding of T2I diffusion models trained on it, while its size is only a fraction of web-scale datasets (0.004% of LAION-400M [schuhmann_laion-400m_2021], or 0.13% of CC-12M [changpinyo_conceptual_2021], showcasing the efficiency and efficacy of the SCOP data engine.
3.2 Token Encoding Ordering Module
Table 1: Most similar prompt variation retrieved by different text encoders.
Text Encoder Used by $F_{\text{text}}$ Space Most similar prompt variation Correct
rephrased negated relation swapped entities
CLIP ViT-L [radford_learning_2021] All 768 1 5088 1231 0.02%
OpenCLIP ViT-H [cherti_reproducible_2023] SD2.1 1024 0 6054 266 0%
OpenCLIP ViT-bigG [cherti_reproducible_2023] SDXL [podell_sdxl_2023], SD3 [esser_scaling_2024], SD3.5 1280 2 6067 251 0.03%
T5-XXL [raffel_exploring_2020] SD3 [esser_scaling_2024], SD3.5, FLUX.1 4096 306 4777 1237 4.84%
Text-to-image diffusion models rely on text encoders to transform natural language descriptions into semantic representations that guide the image generation process. For generation of accurate spatial relationships, these semantic representations must preserve the accurate spatial relationships described in the input text. This is the cornerstone for effective spatial understanding of T2I diffusion models.
Analysis of Spatial Understanding.
We investigate whether the spatial information is properly represented for the diffusion model to understand it. Our key insight is that if text encoders were able to properly preserve spatial relationships, then logically equivalent spatial descriptions should yield high similarities in their encoded representations. A proxy task is designed to quantify this property: given a base prompt describing a spatial relationship (e.g., “A to the left of B”), we generate three variations of it: (1) rephrased: logically equivalent but differently worded (“B to the right of A”), (2) negated relation: substituting the relation phrase with its opposite (“A to the right of B”), and (3) swapped entities: exchanging object positions (“B to the left of A”). Theoretically, only the rephrased variation should produce an encoding similar to the base prompt.
To ensure comprehensive evaluation, we conduct this analysis using all 80 object categories from COCO [lin_microsoft_2014], with four fundamental spatial relationships (“left”, “right”, “above”, “below”), yielding 6,320 prompts. We retrieve the most similar prompt variation according to the similarity between the encoded representations across four text encoders. The results are shown in Tab. 1: even T5-XXL [raffel_exploring_2020] with 11B parameters fails to recognize the logically equivalent variation over 95% of the time. Lighter encoders like CLIP [radford_learning_2021, cherti_reproducible_2023] show near-complete failures across different model sizes, suggesting that current text encoders do not sufficiently preserve spatial relationships in their encoded representations.
Solution.
Figure 4: Overview of Token Encoding Ordering (TENOR). TENOR injects token ordering information into every text-image attention operation in either UNet- or MMDiT-based diffusion models.
Based on this analysis, we propose TENOR, a plug-and-play module designed to compensate for the limitations of current text encoders. TENOR operates by augmenting the conditioning text signals with their original token ordering information in attention operations throughout the text-to-image generation process. Unlike the original positional encoding in transformers which only adds positional encodings to the initial token embeddings [vaswani_attention_2017], TENOR explicitly injects the token ordering information into every text-image attention operation in the diffusion model. This design ensures that spatial semantics are actively preserved every time text guidance influences the image generation process.
Implementation and Efficiency.
Fig. 4 illustrates the overall process of TENOR. Specifically, in every text-image attention operation in the diffusion model, TENOR adds absolute positional encodings [vaswani_attention_2017] to the $K$ vector for UNet-based models [rombach_high-resolution_2022, podell_sdxl_2023], and both $K_{\text{text}}$ and $Q_{\text{text}}$ for MMDiT-based models [peebles_scalable_2023, esser_scaling_2024]. While concurrent approaches have suggested fine-tuning text encoders [chatterjee_getting_2024] during training, performing test-time optimizations [chefer_attend-and-excite_2023, chen_training-free_2023], or incorporating large language models [hu_ella_2024, wu_paragraph--image_2023], such methods add significant computational overhead either during training or at inference time. In contrast, our solution requires no additional trainable parameters beyond the diffusion model itself, and inflicts practically zero computational penalty at inference time. While adapting TENOR requires a brief fine-tuning phase of the diffusion model, this process is significantly more efficient than training new spatial-aware text encoders. Empirically, we find that this approach substantially improves spatial understanding of both UNet-based models and the MMDiT-based FLUX.1.
Figure 5: Qualitative results of models enhanced with CoMPaSS. CoMPaSS enhances the spatial understanding of diffusion models across UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1.
4 Experiments
4.1 Training Details
We validate the effectiveness of CoMPaSS on four popular open-weight text-to-image diffusion models, where SD1.4, SD1.5, and SD2.1 are based on the UNet architecture [rombach_high-resolution_2022], and FLUX.1 is based on the MMDiT architecture [peebles_scalable_2023, esser_scaling_2024]. For UNet-based diffusion models, we train the entire UNet with a learning rate of 5e-6. For MMDiT-based FLUX.1, we train a rank-16 LoRA [hu_lora_2021] with a learning rate of 1e-5 due to limited computational resources. All experiments use an effective batch size of 4. Detailed experimental settings can be found in Appendix A.
4.2 Evaluation Metrics
Table 2: Comparing models enhanced with CoMPaSS and their base models on various spatial-related benchmarks. CoMPaSS sets new state-of-the-arts with substantial gains on every spatial-related benchmark, including VISOR [gokhale_benchmarking_2023], T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (T. Spatial), and Geneval Position [ghosh_geneval_2023] (G. Position).
Method VISOR (%) T. Spatial G. Position
uncond cond 1 2 3 4
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 0.12 0.03
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 0.34 0.46
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 0.08 0.04
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 0.35 0.54
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 0.13 0.07
SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 0.21 0.11
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 0.32 0.51
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 0.18 0.26
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 0.30 0.60
Table 3: Comparison of image fidelity metrics on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024].
[…]
SD1.4 13.20 0.5618
SD1.4 +CoMPaSS 10.67 0.3313
SD1.5 12.82 0.5548
SD1.5 +CoMPaSS 10.89 0.3235
SD2.1 21.65 0.6472
SD2.1 +SPRIGHT [chatterjee_getting_2024] 16.15 0.512
SD2.1 +CoMPaSS 16.96 0.4083
FLUX.1 27.96 0.8737
FLUX.1 +CoMPaSS 26.40 0.6859
We extensively evaluate CoMPaSS on several well-known benchmarks, which are described below:
[…]
  • T2I-CompBench [huang_t2i-compbench_2023-1] and GenEval [ghosh_geneval_2023] are two benchmarks designed for comprehensive evaluation of the text-image alignment capabilities of T2I models. In addition to spatial-related evaluations, they provide methodologies for assessing correctness of attribute binding, non-spatial relationships, and counting. We use them to evaluate the performance impact of CoMPaSS on tasks other than spatial understanding.
  • Fréchet Inception Distance (FID) [heusel_gans_2017] and CLIP Maximum Mean Discrepancy (CMMD)  [jayasumana_rethinking_2024] estimate the fidelity of images generated from T2I models. FID estimates the photorealism of generated images. CMMD is an estimator of image quality that aligns better with human preferences and aims to account for some of the limitations of FID. We report them both.
4.3 Main Results
Spatial Understanding.
We evaluate spatial relationship generation accuracy across multiple benchmarks and report them in Tab. 2. Adding CoMPaSS to existing diffusion models sets new state-of-the-art across all spatial-related metrics. Notably, with the best open-weight diffusion model FLUX.1, CoMPaSS obtains substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (131%). As evident in Fig. 5, while existing models often struggle with spatial relationships, CoMPaSS effectively addresses these limitations even on novel spatial configurations unseen during training, demonstrating significant boosts in spatial understanding.
General Generation Capability and Fidelity.
Apart from spatial-related metrics, we evaluate other tasks in the GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1] benchmarks, and image fidelity scores on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024]. The results are reported in Tabs. 5 and 3. CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores and image fidelity. We conjecture that in the base models, the spatial terms were entangled with unrelated semantics due to severely flawed data and unclear representation. By exclusively disentangling the spatial terms, the model also learns to better understand other aspects of language, resulting in improvements on other tasks.
4.4 Ablation Studies
Table 4: Ablation studies of each component of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Sec. 3.1; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
Setting Model Components T. Spatial G. Pos.
SCOP TENOR
(i) SD1.5 0.08 0.04
(ii) SD1.5 0.32 0.39
(iii) SD1.5 0.35 0.54
(i) FLUX.1 0.18 0.26
(ii) FLUX.1 0.29 0.56
(iii) FLUX.1 0.30 0.60
Table 5: Evaluation results of general generation capabilities across a wide range of tasks on GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1]. CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores.
Method T2I-CompBench GenEval
Color Shape Texture Spatial Non-Spatial Position Single O. Two O. Counting Colors Attr. Overall
SD1.4 0.38 0.36 0.42 0.12 0.31 0.03 0.98 0.41 0.34 0.74 0.06 0.43
SD1.4 +CoMPaSS 0.49 0.43 0.53 0.34 0.31 0.46 0.99 0.68 0.34 0.73 0.17 0.56
SD1.5 0.38 0.37 0.42 0.08 0.31 0.04 0.96 0.38 0.36 0.75 0.06 0.42
SD1.5 +CoMPaSS 0.50 0.43 0.52 0.35 0.31 0.54 0.99 0.69 0.34 0.72 0.15 0.57
SD2.1 0.51 0.42 0.49 0.13 0.31 0.07 0.98 0.51 0.44 0.85 0.17 0.50
SD2.1 +SPRIGHT [chatterjee_getting_2024] - - - 0.21 - 0.11 0.99 0.59 0.49 0.85 0.15 0.51
SD2.1 +CoMPaSS 0.55 0.43 0.54 0.32 0.30 0.51 0.99 0.69 0.20 0.71 0.15 0.54
FLUX.1 0.69 0.48 0.63 0.18 0.31 0.26 0.92 0.77 0.71 0.66 0.27 0.60
FLUX.1 +CoMPaSS 0.83 0.59 0.71 0.30 0.32 0.60 0.99 0.87 0.71 0.80 0.76 0.76
Figure 6: Ablation study on the TENOR module. TENOR helps the model to gain better generalization capability beyond training data.
In the following, we conduct controlled experiments on SD1.5 (most popular and accessible UNet-based diffusion model) and FLUX.1-dev (a state-of-the-art diffusion model based on the MMDiT architecture) to better understand how each component contributes to generation performance. We report the spatial accuracy metrics under various settings in Tab. 4.
Effect of SCOP.
Comparing settings (i) and (ii) within each model reveals that SCOP alone contributes substantially to the spatial understanding of diffusion models. This finding is in line with our initial motivation, where we identified that one of the main reasons current models are failing spatial-related generations is that current datasets yield severely flawed spatial-related image-text data. It also confirms that SCOP is able to curate a set of high-quality data that directly contributes to the spatial understanding of diffusion models.
Effect of TENOR.
Tab. 4 shows that our full method (setting (iii) within each model) further improves the spatial accuracy. We prompt the models with text descriptions unseen during training, and compare the generation results with setting (ii). The results are shown in Fig. 6. While SCOP provides rich and accurate spatial priors, it is only after incorporating TENOR that the model can better generalize to hard unseen spatial configurations. This phenomenon connects closely to our initial finding about the spatial understanding deficiency of popular text encoders in Tab. 1: when the semantic representations of completely opposite text descriptions are indistinguishable, the model is trained to align mixed image signals from seemingly similar text conditions, resulting in poor generalization beyond training data. TENOR compensates this issue by actively injecting token ordering information into every text-image attention operation throughout the diffusion model, disentangling different spatial configurations at the input level, alleviating the model’s burden of distinguishing different spatial terms, as a result achieving better zero-shot generation on unseen settings.
5 Conclusion
We present CoMPaSS, a versatile training framework that enhances spatial understanding of any text-to-image diffusion model. CoMPaSS addresses the problem of ambiguous spatial relationships in datasets through the SCOP data engine, which curates spatially-accurate training data through principled constraints. CoMPaSS additionally uses a parameter-free module, TENOR, that further boosts spatial understanding of T2I diffusion models of any architecture while adding negligible computational overhead. We conduct extensive experiments on four popular open-weight diffusion models, covering both UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1, and set new state-of-the-art across well-known benchmarks with substantial relative gains, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%). Importantly, the enhancement in spatial understanding comes without compromising the general generation capabilities or image quality. By addressing a fundamental limitation of current text-to-image diffusion models, we believe CoMPaSS takes an important step towards reliable spatially-accurate image generation, and opens up new possibilities for applications requiring precise spatial control.
Appendix
Table 6: Comparison to state-of-the-art models on the VISOR [gokhale_benchmarking_2023] benchmark. OA stands for “object accuracy”, which is the success rate of the mentioned objects to appear in the image.
Method uncond cond 1 2 3 4 OA
GLIDE [nichol_glide_2022] 1.98 59.06 6.72 1.02 0.17 0.03 3.36
DALLE-mini [Dayma_DALL·E_Mini_2021] 16.17 59.67 38.31 17.50 6.89 1.96 27.10
CogView2 [ding_cogview2_2022] 12.17 65.89 33.47 11.43 3.22 0.57 18.47
Structured Diffusion [feng_training-free_2023] 17.87 62.36 44.70 18.73 6.57 1.46 28.65
DALLE-2 [ramesh_hierarchical_2022] 37.89 59.27 73.59 47.23 23.26 7.49 63.93
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 29.86
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 28.79
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 47.83
SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 60.68
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 56.95
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 65.56
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 65.78
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 68.23
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 78.64
Appendix A Additional Training Details
Table 7: Hyperparameters used during training.
Hyperparameter SD1.4 SD1.5 SD2.1 FLUX.1
AdamW Learning Rate (LR) 5e-6 5e-6 5e-6 1e-4
AdamW $\beta_{1}$ 0.9 0.9 0.9 0.9
AdamW $\beta_{2}$ 0.999 0.999 0.999 0.999
AdamW $\epsilon$ 1e-8 1e-8 1e-8 1e-8
AdamW Weight Decay 1e-2 1e-2 1e-2 1e-2
LR scheduler Constant Constant Constant Constant
LR warmup steps 0 0 0 20
Training Steps 24,000 24,000 80,000 24,000
Local Batch Size 1 1 1 1
Gradient Accumulation 2 2 2 2
Training GPUs 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S
Training Resolution $512\times 512$ $512\times 512$ $512\times 512$ $512\times 512$
Trained Parameters All parameters of diffusion UNet LoRA (rank=16) on all DoubleStreamBlocks
Prompt Dropout Probability 10% 10% 10% 10%
A.1 SCOP Dataset Construction
For the SCOP data engine, we empirically determined the optimal threshold values $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ through extensive experimentation. We then applied these thresholds to process the COCO training split [lin_microsoft_2014], resulting in the curated SCOP dataset. The initial Relationship Reasoning stage identifies 2,468,858 object pairs with their corresponding spatial relationships. These pairs are then systematically filtered through the Spatial Constraints Enforcement stage, which applies a series of increasingly stringent criteria: Visual Significance (eliminating 1,929,560 pairs), Semantic Distinction (removing 169,973 pairs), Spatial Clarity (filtering out 119,457 pairs), Minimal Overlap (excluding 148,376 pairs), and Size Balance (removing 73,464 pairs), ultimately yielding 28,028 clear and unambiguous object pairs.
During the training phase, the spatial relationships between object pairs is expressed with one of 8 distinct relationship tokens: <left>, <above>, <right>, <below>, <left+above>, <right+above>, <left+below>, and <right+below>. These relationships are determined based on the exact positions of the objects, with occasional random replacements using the <and> token to enhance robustness. For each pair, we select a square region containing both objects, which is then randomly expanded by up to 10% before being cropped to create the final training image paired with its corresponding text description.
A.2 Model Setups
Our experiments are conducted on four diffusion models, including three UNet-based diffusion models SD1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4, SD1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, SD2.1333https://huggingface.co/stabilityai/stable-diffusion-2-1, and a state-of-the-art MMDiT-based diffusion model FLUX.1-dev [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev.
For the UNet-based models, we discovered that incorporating attention supervision as proposed by [wang_tokencompose_2024] significantly enhanced the convergence of the TENOR module. We therefore integrated this supervision across all UNet-based implementations. In contrast, for the FLUX.1 model, we found that the standard denoising loss alone was sufficient to achieve optimal performance. Notably, our best results were achieved with the CoMPaSS-enhanced FLUX.1 model, using a rank-16 LoRA checkpoint that requires only $\sim$50MiB on-disk storage, making it highly efficient for practical applications.
Detailed training hyperparameters for all model configurations are provided in Tab. 7.
Appendix B Runtime Performance of TENOR
The TENOR module represents the only potential source of additional computational overhead in our framework CoMPaSS, as it injects token ordering information into each text-image attention operation within the diffusion models. To quantify its impact, we conducted comprehensive benchmarking of inference latency across all model configurations:
Model Latency @ $512\times 512$ Overhead
SD1.4 1.17s $\pm$ 3.04ms +0.85%
SD1.4 +TENOR 1.18s $\pm$ 7.24ms
SD1.5 1.17s $\pm$ 2.50ms +1.71%
SD1.5 +TENOR 1.19s $\pm$ 4.70ms
SD2.1 1.13s $\pm$ 2.50ms +4.42%
SD2.1 +TENOR 1.18s $\pm$ 4.70ms
FLUX.1 17.3s $\pm$ 40.6ms +2.89%
FLUX.1 +TENOR 17.8s $\pm$ 88.8ms
Our measurements demonstrate that the TENOR module has minimal impact on runtime performance, introducing only negligible computational overhead. Even in the most demanding case of the FLUX.1-dev model, the additional time penalty amounts to just 2.89% of the total inference time, making it a highly practical enhancement for real-world applications.
Appendix C More Experimental Results
C.1 Additional Comparisons on VISOR
We present a comprehensive comparison of VISOR metrics against other state-of-the-art models in Tab. 6, demonstrating the superior performance of our approach across various evaluation criteria.
C.2 More Visual Comparisons
To provide a clear visualization of our approach’s effectiveness, we present additional visual comparisons across eight spatial configurations in Figs. 9, 8, 11 and 10. For each prompt, we generate a total of 36 images using different model configurations, offering a comprehensive view of how our method consistently improves spatial understanding across various scenarios and model architectures.
C.3 Ablation Studies on other Models
Table 8: Ablation studies of each component of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Sec. 3.1 of the main paper; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
Setting Model Components T. Spatial G. Pos.
SCOP TENOR
(i) SD1.4 0.12 0.03
(ii) SD1.4 0.29 0.36
(iii) SD1.4 0.34 0.46
(i) SD2.1 0.13 0.07
(ii) SD2.1 0.30 0.36
(iii) SD2.1 0.32 0.51
While Tab. 4 in the main paper presented ablation studies on SD1.5 and FLUX.1, here we extend our analysis to the other two UNet-based diffusion models SD1.4 and SD2.1 in Tab. 8. The results consistently demonstrate that both components of CoMPaSS contribute to improved spatial understanding across different model architectures.
Figure 7: Example object pairs and their corresponding bounding boxes extracted by the SCOP data engine. Each pair satisfies our spatial constraints for Visual Significance, Semantic Distinction, Spatial Clarity, Minimal Overlap, and Size Balance, ensuring unambiguous spatial relationships.
Figure 8: Additional results demonstrating spatial relationship “left”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 9: Additional results demonstrating spatial relationship “above”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 10: Additional results demonstrating spatial relationship “right”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 11: Additional results demonstrating spatial relationship “below”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.

### v2

@@
 Abstract
-Text-to-image (T2I) diffusion models excel at generating photorealistic images, but commonly struggle to render accurate spatial relationships described in text prompts.
-We identify two core issues underlying this common failure: 1) the ambiguous nature of spatial-related data in existing datasets, and 2) the inability of current text encoders to accurately interpret the spatial semantics of input descriptions.
-We address these issues with CoMPaSS, a versatile training framework that enhances spatial understanding of any T2I diffusion model.
-CoMPaSS solves the ambiguity of spatial-related data with the Spatial Constraints-Oriented Pairing (SCOP) data engine, which curates spatially-accurate training data through a set of principled spatial constraints.
-To better exploit the curated high-quality spatial priors, CoMPaSS further introduces a Token ENcoding ORdering (TENOR) module to allow better exploitation of high-quality spatial priors, effectively compensating for the shortcoming of text encoders.
-Extensive experiments on four popular open-weight T2I diffusion models covering both UNet- and MMDiT-based architectures demonstrate the effectiveness of CoMPaSS by setting new state-of-the-arts with substantial relative gains across well-known benchmarks on spatial relationships generation, including VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%).
-Code will be available at https://github.com/blurgyy/CoMPaSS.
+Text-to-image (T2I) diffusion models excel at generating photorealistic images but often fail to render accurate spatial relationships.
+We identify two core issues underlying this common failure: 1) the ambiguous nature of data concerning spatial relationships in existing datasets, and 2) the inability of current text encoders to accurately interpret the spatial semantics of input descriptions.
+We propose CoMPaSS, a versatile framework that enhances spatial understanding in T2I models.
+It first addresses data ambiguity with the Spatial Constraints-Oriented Pairing (SCOP) data engine, which curates spatially-accurate training data via principled constraints.
+To leverage these priors, CoMPaSS also introduces the Token ENcoding ORdering (TENOR) module, which preserves crucial token ordering information lost by text encoders, thereby reinforcing the prompt’s linguistic structure.
+Extensive experiments on four popular T2I models (UNet and MMDiT-based) show CoMPaSS sets a new state of the art on key spatial benchmarks, with substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%).
+Code is available at github.com/blurgyy/CoMPaSS.
 1 Introduction
-Recent advancements in text-to-image (T2I) diffusion models [saharia_photorealistic_2022, ramesh_hierarchical_2022, rombach_high-resolution_2022] have transformed visual content creation and significantly shaped the modern digital life [betker_improving_nodate, MidJourney, RunwayML].
-These models are capable of synthesizing photorealistic images with remarkable details [dhariwal_diffusion_2021], and are now widely applied across a range of creative and practical tasks [hertz_prompt--prompt_2022, lugmayr_repaint_2022, ruiz_dreambooth_2023, poole_dreamfusion_2022].
-Despite these advancements, however, a key limitation persists: today’s diffusion models frequently fail to correctly render spatial relationships described in text.
-For example, when given seemingly simple spatial configurations like “a motorcycle to the right of a bear”, or “a bird below a skateboard”, models that excel in realism often struggle to accurately capture these spatial relationships (Fig.
-1 left).
-Figure 1: CoMPaSS enables existing T2I diffusion models to generate spatial configurations that are unlikely to occur naturally and unseen during training by improving their spatial understanding.
-Given that T2I diffusion models have shown significant capability in encoding abstract attributes such as aesthetic style and quality [dhariwal_diffusion_2021, rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024], their consistent failure in handling spatial relationships warrants closer examinations.
-We consider that the limitation stems from how spatial priors are organized within the training data, and how these priors are represented to the diffusion model.
-To better understand this limitation, we conduct a thorough investigation covering the data characteristics during training, and the language representation during generation.
+Recent advancements in text-to-image (T2I) diffusion models [62, 57, 58] have transformed visual content creation and significantly shaped the modern digital life [2, 47, 60, 4].
+These models are capable of synthesizing photorealistic images with remarkable details [18], and are now widely applied across a range of creative and practical tasks [30, 45, 59, 54].
+Despite these advancements, a key limitation persists: today’s diffusion models frequently fail to correctly render spatial relationships described in text.
+For example, when given seemingly simple spatial configurations like “a motorcycle to the right of a bear”, or “a bird below a skateboard”, models that excel in realism often struggle to accurately capture these spatial relationships.
+Given that T2I diffusion models have shown significant capability in encoding abstract attributes such as aesthetic style and quality [18, 58, 53, 20, 4], their consistent failure in handling spatial relationships warrants closer examination.
+Figure 1: CoMPaSS enhances the spatial understanding of existing T2I diffusion models, enabling them to generate images that faithfully reflect spatial configurations specified in the text prompt.
+To better understand this limitation, we conduct a thorough investigation covering the essential aspects of training a diffusion model, including the characteristics of the training data and the quality of language representations used during generation.
 Our analysis identified two core issues underlying this spatial understanding deficiency.
-First, the organization of spatial related data in widely-used image-text datasets including LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022], CC-12m [changpinyo_conceptual_2021], and COCO [lin_microsoft_2014] is severely flawed.
+First, the organization of data concerning spatial relationships in widely-used image-text datasets including LAION [64, 65], CC12M [5], and COCO [43] is severely flawed.
 As shown in Fig.
-2, spatial descriptions appear frequently ambiguous due to various reasons, such as inconsistent frame of reference, or non-spatial usage of spatial terms.
+2, spatial descriptions in these datasets are frequently ambiguous due to various reasons, such as inconsistent frames of reference or the non-spatial use of spatial terms.
 This inevitably leads to the spatial terms being interpreted by the diffusion model as ambiguous.
-Second, commonly used text encoders including CLIP [radford_learning_2021, cherti_reproducible_2023] and T5 [raffel_exploring_2020] consistently fail to preserve spatial relationships in their embeddings.
-Through a proxy task where the text encoders are extensively tested to retrieve the spatial equivalent from three prompt variations (Tab.
-1), we show that current text encoders struggle to interpret the accurate meanings of spatial descriptions.
+Second, commonly used text encoders including CLIP [55, 14] and T5 [56] consistently fail to preserve spatial relationships in their embeddings.
+We demonstrate this through a proxy task (Tab.
+1) where we test the ability of text encoders to identify spatially equivalent descriptions among a set of prompt variations.
+Our results show that these encoders struggle to interpret spatial descriptions accurately.
 Furthermore, the two issues compound each other.
@@
 Conversely, even high-quality spatial descriptions would be undermined by the encoder’s inability to interpret them accurately.
-To this end, we seek to enhance the spatial understanding of text-to-image generative diffusion models with a comprehensive solution, which we call CoMPaSS (Comprehensive Method for Positional and Spatial Synthesis).
+To this end, we propose CoMPaSS (Comprehensive Method for Positional and Spatial Synthesis), a comprehensive solution to enhance the spatial understanding of T2I diffusion models.
 Central to our approach is the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions.
 By enforcing strict criteria for visual prominence, positional clarity, and size balance, SCOP identifies and validates spatial relationships between object pairs through carefully designed spatial constraints.
-When applied to the COCO [lin_microsoft_2014] training split, SCOP yields a curated set of over 28,000 object pairs, each coupled with spatially-accurate text descriptions of the image content.
-To better exploit the accurate spatial priors, we additionally introduce Token ENcoding ORdering (TENOR), a plug-and-play module that enhances the preservation of spatial semantics from text input with explicit token ordering information, effectively compensating for the spatial understanding limitations of current text encoders.
-TENOR is compatible with any diffusion architecture, adds no extra parameters, and inflicts practically zero inference-time overhead.
-Applying CoMPaSS to popular open-weight T2I diffusion models like SD1.4, SD1.5, SD2.1 and FLUX.1 sets new state-of-the-arts with substantial relative gains on well-known benchmarks, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%).
+When applied to the COCO [43] training split, SCOP yields a curated set of over 28,000 object pairs, each coupled with spatially-accurate text descriptions.
+To better exploit the spatial priors curated by SCOP, we introduce the Token ENcoding ORdering (TENOR) module.
+This plug-and-play module addresses a key limitation of text encoders by reinforcing the prompt’s linguistic structure.
+It does so by injecting explicit token ordering information into the model’s attention mechanism, ensuring that the sequential nature of the text is preserved.
+TENOR is compatible with any diffusion architecture, adds no extra parameters, and incurs negligible inference-time overhead.
+When applied to popular open-weight T2I diffusion models (SD1.4, SD1.5, SD2.1, and FLUX.1), CoMPaSS achieves state-of-the-art performance on well-known benchmarks, with substantial relative gains on VISOR [26] (+98%), T2I-CompBench Spatial [35] (+67%), and GenEval Position [25] (+131%).
+Notably, these improvements are achieved without compromising general generation capabilities, while also enhancing image fidelity.
 In summary, our contributions are three-fold:
@@
   • A systematic data engine, SCOP, that enforces principled constraints to identify and validate unambiguous spatial relationships between object pairs in images, enabling the curation of high-quality spatial training data.
-  • A parameter-free module, TENOR, that further improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds negligible computational overhead.
-2 Related Works
+  • A parameter-free module, TENOR, that improves a model’s ability to interpret spatial language by preserving the crucial token order of text prompts, while adding negligible computational overhead.
+2 Related Work
 Text-to-image Diffusion Models.
-Diffusion models have achieved very high-quality and diverse image synthesis in recent years [ho_denoising_2020, rombach_high-resolution_2022, dhariwal_diffusion_2021, peebles_scalable_2023].
-To allow natural user input through human language, text-conditional diffusion models adopt datasets with large-scale image-text pairs [changpinyo_conceptual_2021, schuhmann_laion-400m_2021, schuhmann_laion-5b_2022, liang_rich_2024], with pre-trained text encoders to bridge the gap between the two modalities.
-GLIDE [nichol_glide_2022], DALLE-2 [ramesh_hierarchical_2022], LDM [rombach_high-resolution_2022], and the Stable-Diffusion family of models [rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024] employ CLIP [radford_learning_2021, cherti_reproducible_2023] for language conditioning.
-PixArt-$\alpha$ [chen_pixart-alpha_2023], SD3/3.5 [esser_scaling_2024], and FLUX.1 adopt the pure-language T5 [raffel_exploring_2020] model as their text encoders.
-Benchmarking the alignment of image content w.r.t.
-the input text description has been well explored [hu_tifa_2023, ghosh_geneval_2023, gokhale_benchmarking_2023, huang_t2i-compbench_2023-1, jayasumana_rethinking_2024, hu_ella_2024].
-More recently, CoMat [jiang_comat_2024] leverages image-to-text (I2T) models [li_blip_2022, wang_git_2022, liu_visual_2023] to identify the overlooked condition information in generated images during training.
-ParaDiffusion [wu_paragraph--image_2023] and ELLA [hu_ella_2024] explore the use of large language models (LLMs) in T2I diffusion models.
-Instead of trying to enhance the language understanding ability by using more advanced text encoders, we seek an orthogonal direction of facilitating text-image alignment with a data engine that ensures accurate spatial alignment between image and text in training data.
+Diffusion models have achieved very high-quality and diverse image synthesis in recent years [32, 58, 18, 51, 12, 11, 10].
+To enable user input via natural language, text-conditional diffusion models are trained on large-scale image-text datasets [5, 64, 65, 42], typically using pre-trained text encoders to bridge the two modalities.
+While early models such as GLIDE [49], DALLE-2 [57], and the Stable Diffusion family [58, 53, 20] employed CLIP [55, 14] for language conditioning, more recent work like PixArt [12, 11, 10], SD3 [20], and FLUX.1 has adopted text-only encoders like T5 [56] to further enhance faithfulness to the input prompt.
+To quantify this faithfulness, benchmarks with diverse evaluation methodologies have been explored [34, 25, 26, 35, 33, 39, 24, 72, 62, 83].
+Although recent reinforcement learning-based methods [21, 61, 38, 69, 3, 27, 35], which often optimize for human preference using specialized datasets [42, 66, 1, 76, 75, 80, 35], have exhibited strong benchmark scores, the ambiguity underlying these datasets remains a significant challenge.
+Our work addresses this issue by introducing a data engine designed to ensure accurate spatial alignment between images and text in the training data.
 Spatial Control in Diffusion Models.
-Controllable image generation has become a popular research topic recently.
-Several lines of works, including both training-based and test-time optimization methods have been explored to enhance the spatial controllability of generation.
-Previous works add image as conditioning input [zhang_adding_2023, mou_t2i-adapter_2023, ye_ip-adapter_2023] to enable layout control, but add computational overhead to both training and inference processes.
-Attention-based methods build upon the relation between the text-image attention map and the layout of generated image [hertz_prompt--prompt_2022], and improve compositional generation by explicitly supervising the attention maps either during training [wang_tokencompose_2024] or at inference-time [chefer_attend-and-excite_2023, chen_training-free_2023, phung_grounded_2024].
-However, supervising attention requires explicit target signal for optimization.
-Although works like [phung_grounded_2024, feng_layoutgpt_2024] automate the generation of target signals with LLMs, optimization on attention maps still inflicts substantial memory and speed penalty.
-SPRIGHT [chatterjee_getting_2024] avoids test-time optimizations by fine-tuning a diffusion model on datasets recaptioned with a vision-language model [liu_visual_2023], but requires joint training of the text encoder in addition to the diffusion network, which still adds considerable training overhead.
-In contrast to existing works, we propose a network module that improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds no extra parameters and inflicts practically zero computational overheads to either training or inference processes.
+Spatial control is one of the most popular demands in image generation [28, 63, 85, 29, 77, 50].
+Approaches for controllable text-to-image generation can be broadly categorized by whether they manipulate the training or the inference stage of a diffusion model.
+Training-based methods [59, 84, 48, 81, 6, 71] adapt the pretrained weights of diffusion models for specific downstream tasks [78, 40, 70, 82, 73, 71, 6, 33].
+However, they generally incur significant training overhead due to architectural augmentations [84, 33, 74], extra training modules [84, 6], or the need for additional supervision [71, 37].
+Inference-only methods [13, 79, 52, 44, 22, 7, 30, 67, 46] achieve robust targeted generation and editing, but often rely on explicit inputs like bounding boxes.
+These inputs may require manual annotation or specialist models that have separate training costs [23, 15] or high inference costs [41, 79].
+Furthermore, many of these approaches introduce significant inference overhead due to costly test-time optimization.
+In contrast, our work introduces a two-part solution.
+Our SCOP data engine provides high-quality training data, while our TENOR module ensures the model can leverage this data by preserving the crucial token order of text prompts.
+This combined approach efficiently improves spatial understanding without extra trainable parameters or significant computational overhead.
 3 Approach
@@
 In Sec.
-3.1, we develop the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions.
+3.1, we introduce the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions.
 In Sec.
-3.2, we introduce the Token ENcoding ORdering (TENOR) module, a plug-and-play module that enhances the preservation of spatial semantics from text input by incorporating explicit token ordering information.
-These components together form a comprehensive approach to improving spatial understanding in text-to-image diffusion models: SCOP provides high-quality spatial priors through carefully curated spatial relationships, while TENOR enables the model to effectively interpret and exploit these spatial priors during generation.
+3.2, we introduce the Token ENcoding ORdering (TENOR) module, a plug-and-play module that improves the model’s ability to process spatial language by preserving the crucial token order of text prompts.
+These two components work in synergy: SCOP provides high-quality spatial priors through carefully curated spatial relationships, while TENOR ensures the model can distinguish between structurally different prompts, allowing it to better leverage these spatial priors during generation.
 3.1 The SCOP Data Engine
@@
 As shown in Fig.
-2, 1) terms like “left” and “right” suffer from perspective ambiguity, referring inconsistently to viewer perspective or object-intrinsic orientation; 2) directional terms are frequently used in non-spatial contexts (e.g.the right choice”); 3) spatial relationships are often described with missing or incorrect reference objects, making the intended spatial configuration impossible to determine.
+2, 1) terms like “left” and “right” suffer from perspective ambiguity, referring inconsistently to viewer perspective or object-intrinsic orientation; 2) directional terms are frequently used in non-spatial contexts (e.g.
+“the right choice”); 3) spatial relationships are often described with missing or incorrect reference objects, making the intended spatial configuration impossible to determine.
 These issues create significant challenges for models attempting to learn reliable spatial relationships from such data.
@@
 3, SCOP processes images through three key stages:
-((a)) COCO [lin_microsoft_2014]
-((b)) LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022]
-((c)) CC-12M [changpinyo_conceptual_2021]
-Figure 2: Examples highlighting common ambiguities and errors in spatial language annotations from COCO, LAION, and CC-12M datasets.
+((a)) COCO [43]
+((b)) LAION [64, 65]
+((c)) CC12M [5]
+Figure 2: Examples highlighting common ambiguities and errors in spatial language annotations from COCO, LAION, and CC12M datasets.
 (a1, a2) inconsistent frame of reference; (b1) non-spatial usage of spatial terms; (b2, c1, c2) missing or incorrect reference objects.
 Figure 3: Overview of the Spatial Constraints-Oriented Pairing (SCOP) data engine.
-SCOP first (1) reasons through all possible pairs of objects in image, then (2) identifies and validates spatial relationships between object pairs through carefully designed spatial constraints.
-Finally, it (3) decodes the extracted unambiguous object pair descriptors into images with accurate spatial descriptions
-1. Relationship Reasoning.
+SCOP first (1) reasons about all possible object pairs in an image, then (2) validates their spatial relationships using a set of carefully designed constraints.
+Finally, it (3) decodes the resulting unambiguous descriptors into image crops paired with accurate textual descriptions.
+1.
+Relationship Reasoning.
 The engine begins by identifying all possible pairwise relationships within each image.
 For every image $\mathbf{I}$, we first identify all valid object instances $\{O_{1},\cdots,O_{n}\}$, where the $i$-th object $O_{i}$ is associated with a bounding box $\mathbf{B}_{i}$ and its corresponding category label $\text{Category}(\mathbf{B}_{i})$.
-We then enumerate all possible pairs between the identified objects, creating a total of $(^{n}_{2})$ candidate pairs with their respective spatial relationships to evaluate.
-2. Spatial Constraints Enforcement.
+We then enumerate all possible pairs of identified objects, creating a total of $\binom{n}{2}$ candidate pairs to be evaluated for their spatial relationships.
+2.
+Spatial Constraints Enforcement.
 To ensure unambiguous spatial relationships between object pairs in images, we formalize a set of necessary constraints that collectively guarantee visual clarity and semantic validity.
@@
     $\frac{\text{Area}(\mathbf{B}_{i}\cup\mathbf{B}_{j})}{\text{Area}(\mathbf{I})}>\tau_{\text{v}}.$ (1)
-    A relatively small value of $\tau_{\text{v}}$ ensures the reasoned spatial relationship from last stage is a dominant feature of the image, rather than an incidental arrangement of background elements.
-  • Semantic Distinction: Objects must belong to different categories, i.e.:
+    A relatively small value of $\tau_{\text{v}}$ ensures that the identified spatial relationship is a dominant feature of the image, rather than an incidental arrangement of background elements.
+  • Semantic Distinction: Objects must belong to different categories, satisfying the following condition:
     $\text{Category}(\textbf{B}_{i})\neq\text{Category}(\textbf{B}_{j}).$ (2)
@@
   • Spatial Clarity: Objects must be within a reasonable proximity for clear spatial reference.
-Formally, let $d(i,j)=\left\lVert\text{centroid}(\mathbf{B}_{i})-\text{centroid}(\mathbf{B}_{j})\right\rVert_{2}$, and $l(i,j)=\min\left(\text{diagonal}(\mathbf{B}_{i}),\text{diagonal}(\mathbf{B}_{j})\right)$, an object pair $(O_{i},O_{j})$ is said to satisfy spatial clarity when:
+Formally, we define the distance between centroids as $d(i,j)=\left\lVert\text{centroid}(\mathbf{B}_{i})-\text{centroid}(\mathbf{B}_{j})\right\rVert_{2}$ and a characteristic length as $l(i,j)=\min\left(\text{diagonal}(\mathbf{B}_{i}),\text{diagonal}(\mathbf{B}_{j})\right)$.
+An object pair $(O_{i},O_{j})$ satisfies this constraint when:
     $\frac{d(i,j)}{l(i,j)}<\tau_{\text{u}},$ (3)
@@
     $\frac{\min\left(\text{Area}(\mathbf{B}_{i}),\text{Area}(\mathbf{B}_{j})\right)}{\max\left(\text{Area}(\mathbf{B}_{i}),\text{Area}(\mathbf{B}_{j})\right)}>\tau_{\text{s}}.$ (5)
-    With this constraint, both objects should contribute similarly to the spatial relationship, preventing cases where one object is too small to serve as a reliable spatial reference.
+    This constraint ensures that both objects contribute similarly to the spatial relationship, preventing cases where one object is too small to serve as a reliable spatial reference.
 These constraints effectively filter out ambiguous spatial relationships while preserving natural object interactions, creating a strong foundation for clear and consistent spatial descriptions.
-3. Relationship Decoding.
+3.
+Relationship Decoding.
 For object pairs satisfying all constraints, we adopt a representation system that encodes the spatial relationships with structured descriptors, rather than fixed text.
-Such descriptors can be decoded into image-text pairs at training time.
-For example, a descriptor $D=$(cup, $\mathbf{B}_{\mathtt{cup}}$) <above> (couch, $\mathbf{B}_{\mathtt{couch}}$)” can be decoded into $(\mathbf{I}_{\text{roi}},\mathbf{T})$, where $\mathbf{I}_{\text{roi}}$ is a crop of the original image that contains both $\mathbf{B}_{\mathtt{cup}}$ and $\mathbf{B}_{\mathtt{couch}}$, and $\mathbf{T}$ is a text prompt formatted from a predefined pool of templates, e.g.a cup on top of a couch”, or “an image of a couch below a cup”, etc.
-Such a decoding process captures the exact positions of the referred objects as well as their reasoned spatial relation, allowing for both accurate relation descriptions and flexible training signals.
+These descriptors are then decoded into image-text pairs at training time.
+For example, a descriptor such as D=D=(cup, $\mathbf{B}_{\mathtt{cup}}$) <above> (couch, $\mathbf{B}_{\mathtt{couch}}$)” is decoded into a pair $(\mathbf{I}_{\text{roi}},\mathbf{T})$.
+Here, $\mathbf{I}_{\text{roi}}$ is a crop of the original image containing both $\mathbf{B}_{\mathtt{cup}}$ and $\mathbf{B}_{\mathtt{couch}}$, and $\mathbf{T}$ is a text prompt generated from a predefined pool of templates (e.g., “a cup on top of a couch” or “an image of a couch below a cup”).
+This decoding process captures the exact positions of the referenced objects and their identified spatial relationship, providing both accurate textual descriptions and flexible training signals.
 The SCOP Dataset.
-We automate SCOP on the COCO [lin_microsoft_2014] training split to curate a dataset specifically targeted at training T2I models with improved spatial understanding capabilities, covering over 28,000 object pairs across 15,000 images, with well-defined spatial relationships and accurate spatial descriptors.
-This dataset significantly improves the spatial understanding of T2I diffusion models trained on it, while its size is only a fraction of web-scale datasets (0.004% of LAION-400M [schuhmann_laion-400m_2021], or 0.13% of CC-12M [changpinyo_conceptual_2021], showcasing the efficiency and efficacy of the SCOP data engine.
+We apply the SCOP engine to the COCO [43] training split to curate a dataset for improving the spatial understanding of T2I models.
+The resulting dataset contains over 28,000 object pairs from 15,000 images, each characterized by a well-defined spatial relationship and an accurate textual descriptor.
+The symbolic descriptors used during SCOP’s curation process are internal to the data engine and are not used at inference time.
+The final trained model takes only standard natural language prompts as input.
+The small size of this dataset, which is only a fraction of web-scale datasets like LAION-400M (0.004%) [64] and CC12M (0.13%) [5], highlights the data efficiency of our targeted curation approach.
+To validate the quality of our curation process, we conduct a human study following VISOR’s evaluation protocol, finding an 85.2% agreement rate between human annotations and our generated image-text pairs, indicating strong alignment.
 3.2 Token Encoding Ordering Module
-Table 1: Most similar prompt variation retrieved by different text encoders.
-Text Encoder Used by $F_{\text{text}}$ Space Most similar prompt variation Correct
-rephrased negated relation swapped entities
-CLIP ViT-L [radford_learning_2021] All 768 1 5088 1231 0.02%
-OpenCLIP ViT-H [cherti_reproducible_2023] SD2.1 1024 0 6054 266 0%
-OpenCLIP ViT-bigG [cherti_reproducible_2023] SDXL [podell_sdxl_2023], SD3 [esser_scaling_2024], SD3.5 1280 2 6067 251 0.03%
-T5-XXL [raffel_exploring_2020] SD3 [esser_scaling_2024], SD3.5, FLUX.1 4096 306 4777 1237 4.84%
 Text-to-image diffusion models rely on text encoders to transform natural language descriptions into semantic representations that guide the image generation process.
-For generation of accurate spatial relationships, these semantic representations must preserve the accurate spatial relationships described in the input text.
-This is the cornerstone for effective spatial understanding of T2I diffusion models.
+To generate accurate spatial relationships, these representations must faithfully preserve the spatial information described in the input text.
+This capability is a cornerstone for effective spatial understanding in T2I diffusion models.
 Analysis of Spatial Understanding.
-We investigate whether the spatial information is properly represented for the diffusion model to understand it.
+To investigate this potential failure point, we conduct a systematic analysis of how well current text encoders represent spatial information.
 Our key insight is that if text encoders were able to properly preserve spatial relationships, then logically equivalent spatial descriptions should yield high similarities in their encoded representations.
-A proxy task is designed to quantify this property: given a base prompt describing a spatial relationship (e.g., “A to the left of B”), we generate three variations of it: (1) rephrased: logically equivalent but differently worded (“B to the right of A”), (2) negated relation: substituting the relation phrase with its opposite (“A to the right of B”), and (3) swapped entities: exchanging object positions (“B to the left of A”).
-Theoretically, only the rephrased variation should produce an encoding similar to the base prompt.
-To ensure comprehensive evaluation, we conduct this analysis using all 80 object categories from COCO [lin_microsoft_2014], with four fundamental spatial relationships (“left”, “right”, “above”, “below”), yielding 6,320 prompts.
-We retrieve the most similar prompt variation according to the similarity between the encoded representations across four text encoders.
+To quantify this property, we design a proxy task: given a base prompt describing a spatial relationship (e.g., “A to the left of B”), we generate three variations of it: (1) rephrased: logically equivalent but differently worded (“B to the right of A”), (2) negated relation: substituting the relation phrase with its opposite (“A to the right of B”), and (3) swapped entities: exchanging object positions (“B to the left of A”).
+Ideally, an encoder that understands the spatial layout should yield the highest similarity for the rephrased variation.
+To ensure comprehensive evaluation, we conduct this analysis using all 80 object categories from COCO [43], with four fundamental spatial relationships (“left”, “right”, “above”, “below”), yielding 6,320 prompts.
+For each base prompt, we identify which of the three variations yields the most similar embedding, testing this across four different text encoders.
 The results are shown in Tab.
-1: even T5-XXL [raffel_exploring_2020] with 11B parameters fails to recognize the logically equivalent variation over 95% of the time.
-Lighter encoders like CLIP [radford_learning_2021, cherti_reproducible_2023] show near-complete failures across different model sizes, suggesting that current text encoders do not sufficiently preserve spatial relationships in their encoded representations.
+1: even T5-XXL [56] with 11B parameters fails to recognize the logically equivalent variation over 95% of the time.
+Lighter encoders like CLIP [55, 14] show near-complete failures across different model sizes, suggesting that current text encoders do not sufficiently preserve spatial relationships in their encoded representations.
+Table 1: Most similar prompt variation retrieved by different text encoders. Each encoder is used by at least one popular open-weight diffusion model (SD1.5/2.1/3/XL, FLUX.1).
+Text Encoder Most Similar Variation Correct
+Rephrased Neg. Rel. Swp.
+CLIP ViT-L [55] 1 5088 1231 0.02%
+OpenCLIP ViT-H [14] 0 6054 266 0%
+OpenCLIP ViT-bigG [14] 2 6067 251 0.03%
+T5-XXL [56] 306 4777 1237 4.84%
 Solution.
-Figure 4: Overview of Token Encoding Ordering (TENOR). TENOR injects token ordering information into every text-image attention operation in either UNet- or MMDiT-based diffusion models.
-Based on this analysis, we propose TENOR, a plug-and-play module designed to compensate for the limitations of current text encoders.
-TENOR operates by augmenting the conditioning text signals with their original token ordering information in attention operations throughout the text-to-image generation process.
-Unlike the original positional encoding in transformers which only adds positional encodings to the initial token embeddings [vaswani_attention_2017], TENOR explicitly injects the token ordering information into every text-image attention operation in the diffusion model.
-This design ensures that spatial semantics are actively preserved every time text guidance influences the image generation process.
+Based on this analysis, we propose TENOR, a plug-and-play module designed to address the structural ambiguity limitations of current text encoders.
+TENOR operates by augmenting the text conditioning signals with their original token ordering information within the text-image attention layers of the diffusion model.
+Unlike standard positional encodings in transformers, which are added only once to the initial token embeddings [68], TENOR explicitly injects this token ordering information into every text-image attention operation.
+This design ensures that the prompt’s structural and sequential information is consistently available every time text guidance influences the image generation process.
+Figure 4: Overview of Token Encoding Ordering (TENOR). TENOR injects token ordering information into every text-image attention operation in both UNet- and MMDiT-based diffusion models.
 Implementation and Efficiency.
@@
 4 illustrates the overall process of TENOR.
-Specifically, in every text-image attention operation in the diffusion model, TENOR adds absolute positional encodings [vaswani_attention_2017] to the $K$ vector for UNet-based models [rombach_high-resolution_2022, podell_sdxl_2023], and both $K_{\text{text}}$ and $Q_{\text{text}}$ for MMDiT-based models [peebles_scalable_2023, esser_scaling_2024].
-While concurrent approaches have suggested fine-tuning text encoders [chatterjee_getting_2024] during training, performing test-time optimizations [chefer_attend-and-excite_2023, chen_training-free_2023], or incorporating large language models [hu_ella_2024, wu_paragraph--image_2023], such methods add significant computational overhead either during training or at inference time.
-In contrast, our solution requires no additional trainable parameters beyond the diffusion model itself, and inflicts practically zero computational penalty at inference time.
+Specifically, for UNet-based models [58, 53], TENOR adds absolute positional encodings [68] to the key ($K$) vectors in each text-image attention layer.
+For MMDiT-based models [51, 20], it adds these encodings to both the text query ($Q_{\text{text}}$) and key ($K_{\text{text}}$) vectors.
+Alternative approaches for improving spatial control involve fine-tuning the text encoder [6], performing test-time optimization [7, 13], or incorporating large language models [33, 74].
+However, these methods typically introduce significant computational overhead during either training or inference.
+In contrast, our solution requires no additional trainable parameters and incurs negligible computational overhead at inference time.
 While adapting TENOR requires a brief fine-tuning phase of the diffusion model, this process is significantly more efficient than training new spatial-aware text encoders.
 Empirically, we find that this approach substantially improves spatial understanding of both UNet-based models and the MMDiT-based FLUX.1.
-Figure 5: Qualitative results of models enhanced with CoMPaSS. CoMPaSS enhances the spatial understanding of diffusion models across UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1.
+It is crucial to clarify the distinct roles of TENOR and the SCOP dataset.
+TENOR itself is agnostic to the semantic meaning of spatial terms like “left” or “right”.
+Its function is to provide a clear, unambiguous structural signal about the token order that is often lost by standard text encoders (as shown in Tab.
+1).
+By explicitly preserving the sequence, TENOR ensures that prompts like “A left of B” and “B left of A” produce structurally different conditioning signals for the diffusion model.
+During fine-tuning, the model then learns to associate these distinct structural signals with the correct visual outcomes provided by the high-quality, unambiguous image-text pairs from our SCOP dataset.
+This synergy is what enables the significant improvement in spatial understanding: TENOR providing the structural capacity to differentiate prompts, while SCOP provides the semantic ground truth.
+Figure 5: Qualitative results of models enhanced with CoMPaSS.
+Our method improves the spatial understanding of both UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1.
+More results are available in the supplementary material.
 4 Experiments
-4.1 Training Details
-We validate the effectiveness of CoMPaSS on four popular open-weight text-to-image diffusion models, where SD1.4, SD1.5, and SD2.1 are based on the UNet architecture [rombach_high-resolution_2022], and FLUX.1 is based on the MMDiT architecture [peebles_scalable_2023, esser_scaling_2024].
-For UNet-based diffusion models, we train the entire UNet with a learning rate of 5e-6.
-For MMDiT-based FLUX.1, we train a rank-16 LoRA [hu_lora_2021] with a learning rate of 1e-5 due to limited computational resources.
-All experiments use an effective batch size of 4.
-Detailed experimental settings can be found in Appendix A.
-4.2 Evaluation Metrics
-Table 2: Comparing models enhanced with CoMPaSS and their base models on various spatial-related benchmarks.
-CoMPaSS sets new state-of-the-arts with substantial gains on every spatial-related benchmark, including VISOR [gokhale_benchmarking_2023], T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (T.
-Spatial), and Geneval Position [ghosh_geneval_2023] (G.
-Position).
-Method VISOR (%) T. Spatial G. Position
-uncond cond 1 2 3 4
-SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 0.12 0.03
-SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 0.34 0.46
-SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 0.08 0.04
-SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 0.35 0.54
-SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 0.13 0.07
-SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 0.21 0.11
-SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 0.32 0.51
-FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 0.18 0.26
-FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 0.30 0.60
-Table 3: Comparison of image fidelity metrics on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024].
+We validate the effectiveness of CoMPaSS on four popular open-weight text-to-image diffusion models.
+These models include three UNet-based architectures (SD1.4, SD1.5, and SD2.1) [58] and one MMDiT-based architecture (FLUX.1) [4].
+Detailed experimental settings are provided in the supplementary material.
+4.1 Evaluation Details
+Table 2: Comparison of base models with their CoMPaSS-enhanced counterparts.
+Our method sets a new state of the art on all spatial benchmarks, including VISOR [26], T2I-CompBench Spatial [35] (T.
+Spat.), GenEval Position [25] (Pos.), and DPG-Bench Relation [33] (Rel.).
+We also report overall (Ovr.) scores for general compositional ability on GenEval and DPG-Bench, and image fidelity scores using FID [31] and CMMD [36].
+Detailed per-benchmark results are in the supplementary material.
+Method VISOR (%) T2I-CompBench GenEval DPG-Bench Image Fidelity
 […]
-SD1.4 13.20 0.5618
-SD1.4 +CoMPaSS 10.67 0.3313
-SD1.5 12.82 0.5548
-SD1.5 +CoMPaSS 10.89 0.3235
-SD2.1 21.65 0.6472
-SD2.1 +SPRIGHT [chatterjee_getting_2024] 16.15 0.512
-SD2.1 +CoMPaSS 16.96 0.4083
-FLUX.1 27.96 0.8737
-FLUX.1 +CoMPaSS 26.40 0.6859
-We extensively evaluate CoMPaSS on several well-known benchmarks, which are described below:
+SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 0.12 0.32 0.03 0.43 81.04 62.02 13.20 0.5618
+SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 0.34 0.42 0.46 0.56 83.21 66.07 10.67 0.3313
+SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 0.08 0.31 0.04 0.42 73.49 63.18 12.82 0.5548
+SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 0.35 0.42 0.54 0.57 84.10 65.81 10.89 0.3235
+SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 0.13 0.37 0.07 0.50 83.95 65.47 21.65 0.6472
+SD2.1 +SPRIGHT [6] 43.23 71.24 71.78 51.88 33.09 16.15 0.21 - 0.11 0.51 - - 16.15 0.512
+SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 0.32 0.43 0.51 0.54 86.54 69.48 16.96 0.4083
+FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 0.18 0.46 0.26 0.60 92.30 80.63 27.96 0.8737
+FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 0.30 0.55 0.60 0.76 94.12 84.42 26.40 0.6859
+Table 3: Computational efficiency comparison. Overheads exclude bounding box generation for methods that require it, and are reported in the format “Time/VRAM”. Lower is better.
+Method Training Inference
+SPRIGHT [6] +8.1%/+24.5% -
+R&B [79] - +1108.7%/+150.9%
+Attn. Refocusing [52] - +23.8%/+38.0%
+Layout Guidance [13] - +180.9%/+67.4%
+CoMPaSS (ours) +3.9%/+0.7% +2.47%/+0.6%
 […]
-  • T2I-CompBench [huang_t2i-compbench_2023-1] and GenEval [ghosh_geneval_2023] are two benchmarks designed for comprehensive evaluation of the text-image alignment capabilities of T2I models.
-In addition to spatial-related evaluations, they provide methodologies for assessing correctness of attribute binding, non-spatial relationships, and counting.
-We use them to evaluate the performance impact of CoMPaSS on tasks other than spatial understanding.
-  • Fréchet Inception Distance (FID) [heusel_gans_2017] and CLIP Maximum Mean Discrepancy (CMMD)  [jayasumana_rethinking_2024] estimate the fidelity of images generated from T2I models.
-FID estimates the photorealism of generated images.
-CMMD is an estimator of image quality that aligns better with human preferences and aims to account for some of the limitations of FID.
-We report them both.
+Method # Training Images G. Position
+SD2.1 0 0.07
+SD2.1 +SPRIGHT [6] $\sim$500 0.11
+[…]
+SD2.1 +CoMPaSS 28k 0.51
+FLUX.1 0 0.26
+[…]
+FLUX.1 +CoMPaSS 28k 0.60
+Baselines.
+We compare CoMPaSS against several baselines: 1) SPRIGHT [6] is a training-based method whose focus on spatial consistency is highly aligned with ours.
+2) R&B [79], 3) Attention Refocusing [52], and 4) Layout Guidance [13] are inference-only methods for zero-shot layout control.
+These methods rely on additional image-space guidance, such as bounding boxes or masks, which are typically provided either through manual annotation or by external language models [23, 41, 15].
+Metrics.
+We extensively evaluate CoMPaSS on several benchmarks with varying complexity: VISOR [26] is a specialized benchmark with 31,680 prompts focusing on spatial understanding.
+T2I-CompBench [35] and GenEval [25] are two benchmarks for evaluating open-world compositional text-to-image generation.
+DPG-Bench [33] is a comprehensive dataset consisting of 1,065 lengthy, dense prompts, designed to assess the intricate semantic alignment capabilities of text-to-image models.
+FID [31] and CLIP Maximum Mean Discrepancy (CMMD) [36] assess image fidelity, with FID focusing on photorealism and CMMD providing an alternative measure that better aligns with human preferences.
+Additionally, we measure time and VRAM overhead to evaluate the computational efficiency of each method.
 4.2 Main Results
@@
 2.
-Adding CoMPaSS to existing diffusion models sets new state-of-the-art across all spatial-related metrics.
-Notably, with the best open-weight diffusion model FLUX.1, CoMPaSS obtains substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (131%).
-As evident in Fig.
-5, while existing models often struggle with spatial relationships, CoMPaSS effectively addresses these limitations even on novel spatial configurations unseen during training, demonstrating significant boosts in spatial understanding.
+Adding CoMPaSS to existing diffusion models sets a new state of the art across all spatial-related metrics.
+Notably, when applied to the state-of-the-art open-weight model FLUX.1, CoMPaSS achieves substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), GenEval Position (131%), and the best scores on the DPG-Bench Relation benchmark.
+As shown in Fig.
+5, CoMPaSS effectively addresses these limitations, correctly rendering novel spatial configurations unseen during training and demonstrating a significant enhancement in spatial understanding.
 General Generation Capability and Fidelity.
-Apart from spatial-related metrics, we evaluate other tasks in the GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1] benchmarks, and image fidelity scores on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024].
-The results are reported in Tabs.
-5 and 3.
-CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores and image fidelity.
-We conjecture that in the base models, the spatial terms were entangled with unrelated semantics due to severely flawed data and unclear representation.
-By exclusively disentangling the spatial terms, the model also learns to better understand other aspects of language, resulting in improvements on other tasks.
+In addition to spatial metrics, we evaluate performance on other compositional tasks from the T2I-CompBench [35], GenEval [25], and DPG-Bench [33] benchmarks.
+We also assess image fidelity using FID [31] and CMMD [36].
+The overall results are reported in Tab.
+2, with detailed per-task results provided in Tab.
+A11 in the supplementary material.
+While CoMPaSS is designed to target spatial performance, we find that it also improves overall alignment scores and image fidelity.
+We conjecture that in base models, spatial terms are often entangled with unrelated semantics due to flawed data.
+By disentangling these terms, CoMPaSS may also help the model better understand other aspects of language, resulting in these broader improvements.
+Importantly, on DPG-Bench, which contains 1,065 long natural language prompts, models enhanced with CoMPaSS achieve consistently higher overall scores, indicating that the enhanced spatial understanding generalizes well to longer and more complex text prompts.
+Computational Efficiency.
+We compare the time and VRAM overhead of CoMPaSS to prior work during both training and inference, with results reported in Tab.
+3.
+SPRIGHT [6], a prior training-based method focused on spatial consistency, finetunes the text encoder along with the diffusion network.
+However, this approach inevitably adds computational requirements (+8.1% in time and +24.5% in VRAM).
+Inference-only methods [79, 52, 13] typically enable gradient-based optimization at test time, incurring substantial overhead (e.g.
+time overhead ranges from 23.8% up to 1108.7%).
+In contrast, CoMPaSS only alters the cross-attention calculation and introduces no extra trainable parameters.
+It is therefore lightweight, incurring only $\sim$3% time overhead and less than 1% VRAM overhead during inference.
+[…]
 4.3 Ablation Studies
-Table 4: Ablation studies of each component of CoMPaSS.
-(i) original models; (ii) trained with the SCOP dataset described in Sec.
-3.1; (iii) our full method.
+Table 5: Ablation study on the hyperparameters of SCOP.
+Our main experiments use the hyperparameter set $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$, denoted as “Ours” in the table.
+The results show that performance is robust to variations in these hyperparameters.
+Name $\tau_{\text{v}}$ $\tau_{\text{u}}$ $\tau_{\text{o}}$ $\tau_{\text{s}}$ Ours
+Value 0.1 0.3 1.0 3.0 0.2 0.5 0.3 0.7
+SD1.5 0.48 0.52 0.50 0.49 0.47 0.51 0.51 0.53 0.54
+FLUX.1 0.55 0.56 0.59 0.54 0.54 0.58 0.58 0.59 0.60
+Table 6: Ablation study on the components of CoMPaSS.
+(i) original models; (ii) trained with the SCOP dataset described in Fig.
+3; (iii) our full method.
 T2I-CompBench Spatial (T.
@@
 (iii) FLUX.1 0.30 0.60
-Table 5: Evaluation results of general generation capabilities across a wide range of tasks on GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1].
-CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores.
-Method T2I-CompBench GenEval
-Color Shape Texture Spatial Non-Spatial Position Single O. Two O. Counting Colors Attr. Overall
-SD1.4 0.38 0.36 0.42 0.12 0.31 0.03 0.98 0.41 0.34 0.74 0.06 0.43
-SD1.4 +CoMPaSS 0.49 0.43 0.53 0.34 0.31 0.46 0.99 0.68 0.34 0.73 0.17 0.56
-SD1.5 0.38 0.37 0.42 0.08 0.31 0.04 0.96 0.38 0.36 0.75 0.06 0.42
-SD1.5 +CoMPaSS 0.50 0.43 0.52 0.35 0.31 0.54 0.99 0.69 0.34 0.72 0.15 0.57
-SD2.1 0.51 0.42 0.49 0.13 0.31 0.07 0.98 0.51 0.44 0.85 0.17 0.50
-SD2.1 +SPRIGHT [chatterjee_getting_2024] - - - 0.21 - 0.11 0.99 0.59 0.49 0.85 0.15 0.51
-SD2.1 +CoMPaSS 0.55 0.43 0.54 0.32 0.30 0.51 0.99 0.69 0.20 0.71 0.15 0.54
-FLUX.1 0.69 0.48 0.63 0.18 0.31 0.26 0.92 0.77 0.71 0.66 0.27 0.60
-FLUX.1 +CoMPaSS 0.83 0.59 0.71 0.30 0.32 0.60 0.99 0.87 0.71 0.80 0.76 0.76
-Figure 6: Ablation study on the TENOR module. TENOR helps the model to gain better generalization capability beyond training data.
-In the following, we conduct controlled experiments on SD1.5 (most popular and accessible UNet-based diffusion model) and FLUX.1-dev (a state-of-the-art diffusion model based on the MMDiT architecture) to better understand how each component contributes to generation performance.
-We report the spatial accuracy metrics under various settings in Tab.
-4.
+Figure 6: Qualitative ablation study on the TENOR module.
+The TENOR module improves generalization to unseen prompts.
+(i) First row: original models.
+(ii) Second row: models trained only with the SCOP dataset.
+(iii) Third row: our full method.
+The left three columns show results from SD1.5; the right three columns show results from FLUX.1.
+To understand the contribution of each component, we conduct controlled experiments on two representative models: SD1.5, a popular UNet-based model, and FLUX.1-dev, a state-of-the-art MMDiT-based model.
+Hyperparameters of SCOP.
+The SCOP data engine has four tunable hyperparameters.
+We empirically determine the optimal values to be $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ via grid search.
+In Tab.
+5, we report the model’s sensitivity to each of these four hyperparameters by evaluating performance on nearby values.
+While our chosen hyperparameters yield optimal results, the model’s performance remains high across a range of nearby values, indicating the robustness of the SCOP engine.
 Effect of SCOP.
-Comparing settings (i) and (ii) within each model reveals that SCOP alone contributes substantially to the spatial understanding of diffusion models.
-This finding is in line with our initial motivation, where we identified that one of the main reasons current models are failing spatial-related generations is that current datasets yield severely flawed spatial-related image-text data.
-It also confirms that SCOP is able to curate a set of high-quality data that directly contributes to the spatial understanding of diffusion models.
+In Tab.
+6, comparing settings (i) and (ii) within each model reveals that SCOP alone contributes substantially to the spatial understanding of diffusion models.
+This finding supports our initial hypothesis that a primary cause of failure in spatial generation is the flawed nature of existing training data.
+This confirms that SCOP successfully curates high-quality data that directly improves the spatial understanding of diffusion models.
 Effect of TENOR.
@@
 6 shows that our full method (setting (iii) within each model) further improves the spatial accuracy.
-We prompt the models with text descriptions unseen during training, and compare the generation results with setting (ii).
-The results are shown in Fig.
-6.
-While SCOP provides rich and accurate spatial priors, it is only after incorporating TENOR that the model can better generalize to hard unseen spatial configurations.
-This phenomenon connects closely to our initial finding about the spatial understanding deficiency of popular text encoders in Tab.
-1: when the semantic representations of completely opposite text descriptions are indistinguishable, the model is trained to align mixed image signals from seemingly similar text conditions, resulting in poor generalization beyond training data.
-TENOR compensates this issue by actively injecting token ordering information into every text-image attention operation throughout the diffusion model, disentangling different spatial configurations at the input level, alleviating the model’s burden of distinguishing different spatial terms, as a result achieving better zero-shot generation on unseen settings.
-5 Conclusion
-We present CoMPaSS, a versatile training framework that enhances spatial understanding of any text-to-image diffusion model.
-CoMPaSS addresses the problem of ambiguous spatial relationships in datasets through the SCOP data engine, which curates spatially-accurate training data through principled constraints.
-CoMPaSS additionally uses a parameter-free module, TENOR, that further boosts spatial understanding of T2I diffusion models of any architecture while adding negligible computational overhead.
-We conduct extensive experiments on four popular open-weight diffusion models, covering both UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1, and set new state-of-the-art across well-known benchmarks with substantial relative gains, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%).
-Importantly, the enhancement in spatial understanding comes without compromising the general generation capabilities or image quality.
-By addressing a fundamental limitation of current text-to-image diffusion models, we believe CoMPaSS takes an important step towards reliable spatially-accurate image generation, and opens up new possibilities for applications requiring precise spatial control.
+As shown in Fig.
+6, we evaluate the models on text descriptions unseen during training across all three settings.
+While SCOP provides rich and accurate spatial priors, it is only after incorporating TENOR that the model generalizes better to challenging, unseen spatial configurations.
+This phenomenon connects closely to our initial finding about spatial understanding deficiency in text encoders (Tab.
+1): when the semantic representations of completely opposite text descriptions are indistinguishable, the model is trained to align mixed image signals from seemingly similar text conditions, resulting in poor generalization beyond training data.
+TENOR addresses this issue by injecting token ordering information into the text-image attention operations.
+This ensures that structurally different prompts produce distinct conditioning signals, enabling the model to learn correct associations from the SCOP data and thereby achieve better zero-shot generation on unseen prompts.
+5 Discussion and Conclusion
+Figure 7: Results for “a tiny phone above a large couch”.
+Left: FLUX.1.
+Middle: FLUX.1+CoMPaSS.
+Right: Fine-tuning (100 steps, batch size 4) on FLUX.1+CoMPaSS with only size-contrast data filtered from adapted SCOP.
+Limitations and Future Work.
+While CoMPaSS demonstrates robust performance, its understanding of spatial language is not exhaustive.
+Our framework primarily addresses ambiguity arising from noisy web data and encoder limitations, but does not yet model the rich, context-dependent nature of spatial language, a key topic in Qualitative Spatial Relations (QSR) [9, 8, 16].
+For instance, rendering subjects with extreme size disparities remains a challenge (see Fig.
+7, left and middle), as the interpretation of “above” or “beside” depends heavily on object context.
+Our preliminary experiments show that a targeted fine-tuning stage can mitigate this specific issue (see Fig.
+7, right).
+Furthermore, our current approach is viewer-centric and does not handle object-centric frames (e.g., “to the child’s right hand side”).
+Fully integrating QSR principles and modeling both object-centric and context-dependent relations are important directions for future work.
+Conclusion.
+We have presented CoMPaSS, a versatile framework that enhances spatial understanding in T2I models.
+CoMPaSS uses the SCOP engine to curate high-quality, spatially-unambiguous training data, and the TENOR module to preserve crucial token ordering information that text encoders often lose.
+Extensive experiments show that CoMPaSS sets a new state of the art on key spatial benchmarks, achieving substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%) without compromising general generation capabilities or image quality.
+By addressing a fundamental limitation in T2I models, CoMPaSS represents an important step towards reliable, spatially-aware image generation.
+Acknowledgements
+This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62032011 and the State Key Laboratory of CAD&CG at Zhejiang University.
 Appendix
-Table 6: Comparison to state-of-the-art models on the VISOR [gokhale_benchmarking_2023] benchmark.
-OA stands for “object accuracy”, which is the success rate of the mentioned objects to appear in the image.
-Method uncond cond 1 2 3 4 OA
-GLIDE [nichol_glide_2022] 1.98 59.06 6.72 1.02 0.17 0.03 3.36
-DALLE-mini [Dayma_DALL·E_Mini_2021] 16.17 59.67 38.31 17.50 6.89 1.96 27.10
-CogView2 [ding_cogview2_2022] 12.17 65.89 33.47 11.43 3.22 0.57 18.47
-Structured Diffusion [feng_training-free_2023] 17.87 62.36 44.70 18.73 6.57 1.46 28.65
-DALLE-2 [ramesh_hierarchical_2022] 37.89 59.27 73.59 47.23 23.26 7.49 63.93
-SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 29.86
-SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 28.79
-SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 47.83
-SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 60.68
-FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 56.95
-SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 65.56
-SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 65.78
-SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 68.23
-FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 78.64
 Appendix A Additional Training Details
-Table 7: Hyperparameters used during training.
+A.1 SCOP Dataset Construction
+The optimal threshold values for the SCOP data engine were empirically determined to be $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ through grid search.
+We then applied these thresholds to process the COCO training split [43], resulting in the curated SCOP dataset.
+The initial Relationship Reasoning stage identifies 2,468,858 object pairs with their corresponding spatial relationships.
+The subsequent Spatial Constraints Enforcement stage systematically filters these pairs.
+This process applies a series of increasingly stringent criteria: Visual Significance eliminates 1,929,560 pairs, Semantic Distinction removes 169,973, Spatial Clarity filters out 119,457, Minimal Overlap excludes 148,376, and Size Balance removes a final 73,464 pairs.
+This cascade ultimately yields 28,028 clear and unambiguous object pairs.
+During the training phase, the spatial relationships between object pairs are expressed with one of 8 distinct relationship tokens: <left>, <above>, <right>, <below>, <left+above>, <right+above>, <left+below>, and <right+below>.
+These relationships are determined based on the exact positions of the objects, with occasional random replacements using the <and> token to enhance robustness.
+For each pair, a square region containing both objects is automatically chosen, which is then randomly perturbed and expanded by up to 10% before being cropped to create the final training image paired with its corresponding text description.
+A.2 Model Setups
+Table A7: Hyperparameters used during training.
 Hyperparameter SD1.4 SD1.5 SD2.1 FLUX.1
@@
 Gradient Accumulation 2 2 2 2
-Training GPUs 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S
-Training Resolution $512\times 512$ $512\times 512$ $512\times 512$ $512\times 512$
+[…]
+Training Resolution $\times 512 512×512$$\times 512$ $\times 512 512×512$$\times 512$
 Trained Parameters All parameters of diffusion UNet LoRA (rank=16) on all DoubleStreamBlocks
 Prompt Dropout Probability 10% 10% 10% 10%
-A.1 SCOP Dataset Construction
-For the SCOP data engine, we empirically determined the optimal threshold values $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ through extensive experimentation.
-We then applied these thresholds to process the COCO training split [lin_microsoft_2014], resulting in the curated SCOP dataset.
-The initial Relationship Reasoning stage identifies 2,468,858 object pairs with their corresponding spatial relationships.
-These pairs are then systematically filtered through the Spatial Constraints Enforcement stage, which applies a series of increasingly stringent criteria: Visual Significance (eliminating 1,929,560 pairs), Semantic Distinction (removing 169,973 pairs), Spatial Clarity (filtering out 119,457 pairs), Minimal Overlap (excluding 148,376 pairs), and Size Balance (removing 73,464 pairs), ultimately yielding 28,028 clear and unambiguous object pairs.
-During the training phase, the spatial relationships between object pairs is expressed with one of 8 distinct relationship tokens: <left>, <above>, <right>, <below>, <left+above>, <right+above>, <left+below>, and <right+below>.
-These relationships are determined based on the exact positions of the objects, with occasional random replacements using the <and> token to enhance robustness.
-For each pair, we select a square region containing both objects, which is then randomly expanded by up to 10% before being cropped to create the final training image paired with its corresponding text description.
-A.2 Model Setups
-Our experiments are conducted on four diffusion models, including three UNet-based diffusion models SD1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4, SD1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, SD2.1333https://huggingface.co/stabilityai/stable-diffusion-2-1, and a state-of-the-art MMDiT-based diffusion model FLUX.1-dev [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev.
-For the UNet-based models, we discovered that incorporating attention supervision as proposed by [wang_tokencompose_2024] significantly enhanced the convergence of the TENOR module.
+Our experiments are conducted on four diffusion models, including three UNet-based diffusion models SD1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4, SD1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, SD2.1333https://huggingface.co/stabilityai/stable-diffusion-2-1, and the state-of-the-art MMDiT-based diffusion model FLUX.1-dev [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev.
+For the UNet-based models, we found that incorporating attention supervision as proposed by [71] significantly enhanced the convergence of the TENOR module.
 We therefore integrated this supervision across all UNet-based implementations.
-In contrast, for the FLUX.1 model, we found that the standard denoising loss alone was sufficient to achieve optimal performance.
+In contrast, for the FLUX.1 model, the standard denoising loss alone was sufficient to achieve optimal performance.
 Notably, our best results were achieved with the CoMPaSS-enhanced FLUX.1 model, using a rank-16 LoRA checkpoint that requires only $\sim$50MiB on-disk storage, making it highly efficient for practical applications.
-Detailed training hyperparameters for all model configurations are provided in Tab. 7.
+Detailed training hyperparameters for all model configurations are provided in Tab. A7.
 Appendix B Runtime Performance of TENOR
@@
 Even in the most demanding case of the FLUX.1-dev model, the additional time penalty amounts to just 2.89% of the total inference time, making it a highly practical enhancement for real-world applications.
-Appendix C More Experimental Results
-C.1 Additional Comparisons on VISOR
-We present a comprehensive comparison of VISOR metrics against other state-of-the-art models in Tab. 6, demonstrating the superior performance of our approach across various evaluation criteria.
-C.2 More Visual Comparisons
+Appendix C Additional Results
+C.1 Visualization of SCOP Data
+We provide visualizations of data curated by SCOP in Fig. A8.
+Figure A8: Example object pairs and their corresponding bounding boxes extracted by the SCOP data engine.
+Each pair satisfies our spatial constraints for Visual Significance, Semantic Distinction, Spatial Clarity, Minimal Overlap, and Size Balance, ensuring unambiguous spatial relationships.
+C.2 Additional Comparisons on VISOR
+We present a comprehensive comparison of VISOR metrics against other state-of-the-art models in Tab. A10, demonstrating the superior performance of our approach across various evaluation criteria.
+C.3 More Visual Comparisons
+Figure A9: Additional results demonstrating spatial relationship “left”. Our method consistently improves spatial accuracy over the baseline models.
+Figure A10: Additional results demonstrating spatial relationship “above”. Our method consistently improves spatial accuracy over the baseline models.
+Figure A11: Additional results demonstrating spatial relationship “right”. Our method consistently improves spatial accuracy over the baseline models.
+Figure A12: Additional results demonstrating spatial relationship “below”. Our method consistently improves spatial accuracy over the baseline models.
 To provide a clear visualization of our approach’s effectiveness, we present additional visual comparisons across eight spatial configurations in Figs.
-9, 8, 11 and 10.
+A10, A9, A12 and A11.
 For each prompt, we generate a total of 36 images using different model configurations, offering a comprehensive view of how our method consistently improves spatial understanding across various scenarios and model architectures.
-C.3 Ablation Studies on other Models
-Table 8: Ablation studies of each component of CoMPaSS.
+C.4 Evaluation on Three-Object Compositions
+To further test the generalization capabilities of CoMPaSS, we constructed a more challenging benchmark focused on three-object spatial configurations.
+This benchmark contains 512 prompts, such as “a clock above a broccoli, with a car to the left of the clock”, and is evaluated using an extension of the GenEval methodology [25].
+Despite not being fine-tuned on such complex prompts, CoMPaSS demonstrates a significant improvement in accuracy, as detailed in Tab.
+A8.
+This result builds upon the strong two-object performance shown in the main paper (Tab.
+2).
+Table A8: Multi-object spatial relationship evaluation. Despite being trained only on object pairs, CoMPaSS improves the model’s spatial accuracy on prompts involving three objects.
+Model % Correct (any) % Correct (all)
+SD2.1 9.1 3.3
+SD2.1 +CoMPaSS 32.3 20.8
+FLUX.1 30.12 13.1
+FLUX.1 +CoMPaSS 52.44 31.2
+C.5 Detailed Benchmark Breakdown
+While the main paper reported overall scores for standard benchmarks (Tab.
+2), here we provide a more granular breakdown.
+Table A11 presents the full, per-task results for T2I-CompBench, GenEval [25], and DPG-Bench [33], offering a comprehensive performance overview.
+C.6 Ablation Studies on other Models
+While Tab.
+6 in the main paper presented ablation studies on SD1.5 and FLUX.1, here we extend our analysis to the other two UNet-based diffusion models SD1.4 and SD2.1 in Tab.
+A9.
+The results consistently demonstrate that both components of CoMPaSS contribute to improved spatial understanding across different model architectures.
+Table A9: Ablation study on the components of CoMPaSS.
 (i) original models; (ii) trained with the SCOP dataset described in Sec.
@@
 (iii) SD2.1 0.32 0.51
-While Tab.
-4 in the main paper presented ablation studies on SD1.5 and FLUX.1, here we extend our analysis to the other two UNet-based diffusion models SD1.4 and SD2.1 in Tab.
-8.
-The results consistently demonstrate that both components of CoMPaSS contribute to improved spatial understanding across different model architectures.
-Figure 7: Example object pairs and their corresponding bounding boxes extracted by the SCOP data engine.
-Each pair satisfies our spatial constraints for Visual Significance, Semantic Distinction, Spatial Clarity, Minimal Overlap, and Size Balance, ensuring unambiguous spatial relationships.
-Figure 8: Additional results demonstrating spatial relationship “left”.
-Each row shows generations from different models for the same text prompt.
-The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
-Figure 9: Additional results demonstrating spatial relationship “above”.
-Each row shows generations from different models for the same text prompt.
-The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
-Figure 10: Additional results demonstrating spatial relationship “right”.
-Each row shows generations from different models for the same text prompt.
-The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
-Figure 11: Additional results demonstrating spatial relationship “below”.
-Each row shows generations from different models for the same text prompt.
-The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
+Table A10: Comparison to state-of-the-art models on the VISOR [26] benchmark. OA stands for “object accuracy”, which measures the rate at which all prompted objects appear in the generated image.
+Method uncond cond 1 2 3 4 OA
+GLIDE [49] 1.98 59.06 6.72 1.02 0.17 0.03 3.36
+DALLE-mini [17] 16.17 59.67 38.31 17.50 6.89 1.96 27.10
+CogView2 [19] 12.17 65.89 33.47 11.43 3.22 0.57 18.47
+Structured Diffusion [22] 17.87 62.36 44.70 18.73 6.57 1.46 28.65
+DALLE-2 [57] 37.89 59.27 73.59 47.23 23.26 7.49 63.93
+SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 29.86
+SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 28.79
+SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 47.83
+SD2.1 +SPRIGHT [6] 43.23 71.24 71.78 51.88 33.09 16.15 60.68
+FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 56.95
+SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 65.56
+SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 65.78
+SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 68.23
+FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 78.64
+Table A11: Evaluation results of general generation capabilities across a wide range of tasks on GenEval [25], T2I-CompBench [35], and DPG-Bench [33].
+While designed to target spatial performance, CoMPaSS also improves overall (Ovr.) alignment scores across most tasks.
+Method T2I-CompBench GenEval DPG-Bench
+Spat. Col. Shp. Tex. N.Sp. Pos. S.O. T.O. Count Col. Attr. Ovr. Rel. Glb. Ent. Attr. Other Ovr.
+SD1.4 0.12 0.38 0.36 0.42 0.31 0.03 0.98 0.41 0.34 0.74 0.06 0.43 81.04 78.12 72.23 72.56 59.60 62.02
+SD1.4 +CoMPaSS 0.34 0.49 0.43 0.53 0.31 0.46 0.99 0.68 0.34 0.73 0.17 0.56 83.21 79.33 75.33 72.09 68.00 66.07
+SD1.5 0.08 0.38 0.37 0.42 0.31 0.04 0.96 0.38 0.36 0.75 0.06 0.42 73.49 74.63 74.23 75.39 67.81 63.18
+SD1.5 +CoMPaSS 0.35 0.50 0.43 0.52 0.31 0.54 0.99 0.69 0.34 0.72 0.15 0.57 84.10 82.67 75.20 73.58 60.80 65.81
+SD2.1 0.13 0.51 0.42 0.49 0.31 0.07 0.98 0.51 0.44 0.85 0.17 0.50 83.95 81.16 74.47 75.29 53.60 65.47
+SD2.1 +SPRIGHT [6] 0.21 - - - - 0.11 0.99 0.59 0.49 0.85 0.15 0.51 - - - - - -
+SD2.1 +CoMPaSS 0.32 0.55 0.43 0.54 0.30 0.51 0.99 0.69 0.20 0.71 0.15 0.54 86.54 79.94 78.89 75.39 62.80 69.48
+FLUX.1 0.18 0.69 0.48 0.63 0.31 0.26 0.92 0.77 0.71 0.66 0.27 0.60 92.30 80.55 87.74 85.55 78.40 80.63
+FLUX.1 +CoMPaSS 0.30 0.83 0.59 0.71 0.32 0.60 0.99 0.87 0.71 0.80 0.76 0.76 94.12 82.98 90.53 88.30 82.80 84.42